i have to find the zeros of the function:
f(x)=x^5+2x^4-13x^3-26x^2+36x+72
help me please ):

Answer by lwsshak3(11628)

find the zeros of the function:
f(x)=x^5+2x^4-13x^3-26x^2+36x+72
***
Use Rational Roots Theorem:
...0..|...1......2.....-13.....-26......36.......72
...1..|...1......3.....-10.....-36.......0........72
...2..|...1......4......-5......-36....-36........0 (2 is a root)
...3..|...1......7......16........12......0 (3 is also a root)
==========================
...0..|...1......7......16........12.....
.-1..|...1......6......10.........2.....
.-2..|...1......5.......6..........0 (-2 is a root)
f(x)=(x-2)(x-3)(x+2)(x^2+5x+6)
f(x)=(x-2)(x-3)(x+2)(x+2)(x+3)
zeros are: 2, 3, -2, -2, -3